Suppose that a rectangle has a perimeter of 20 meters. Express the area A(x) of the rectangle in terms of the length x of one of its sides.
Answer by ankor@dixie-net.com(22740)

Suppose that a rectangle has a perimeter of 20 meters.
Express the area A(x) of the rectangle in terms of the length x of one of its sides.
:
Let w = the width, x = the length
:
2x + 2w = 20
Simplify, divide by 2
x + w = 10
w = (10-x)
:
Area
A = x*w
replace w with (10-x)
A(x) = x(10-x)
A(x) = -x^2 + 10x